Solve the system by addition.
x + y = 6
x – y = 4
no book, thanks so much for all you do!!!

Answer by funmath(2933)

Solve the system by addition.
x + y = 6
x – y = 4
_________ Add the two equations together and the y's are eliminated.
2x+0=10
2x=10
2x/2=10/2
x=5 <---- substitute that in for x in either equation:
(5)+y=6
-5+5+y=-5+6
0+y=1
y=1
The solution is: (x,y)=(5,1)